DIRC Operations Weekly Report
Period: Wednesday 2007/03/06 - Wednesday 2007/03/13
Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Bradley Wogsland
File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070313.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070221
PEP peak lumi: ~8.1 x 1033
(previous week: ~8.3 x 1033)
BaBar integrated lumi: ~1.7
(previous week: ~3.1 fb-1)
Run 6 PEP delivery:~1.7
(previous week: ~3.3 fb-1)
Run 6 total delivery: 14.9 fb-1
Run 6 total BaBar recorded: 14.4 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
On March 13 (day shift) some MD studies.
Foreseen work opportunities (PEP down periods):
3 days ROD begining on March 26. Backward door are going to be close, no expected DIRC work.
BaBar activities/issues this week:
- On March 13 at 3:30am DIRC comissioner paged because one scalar (sector 1) went red and kept reading at frequencies of more than 600kHz for about 40 minutes. All the others scalars were fine. That happened during an ODF-CMP problem and it is
maybe related to that.
logbook for details
- Test of DCC and DFBs that arrived from France last week.
- Maybe replace two DFBs in sector 1 with a group of 8 innefficient or flaky channels (require bridge).
- A few noisy channels need to be masked.
- Unmask all masked channels to find which of these are actually noisy. After December DFB work/movement masked channels may no longer be the noisy ones.
Bradley, Jose and Alejandro rotating backup & on-call commissioning duties